### v1

# Speech-Language Models with Decoupled Tokenizers and Multi-Token Prediction

Abstract
Speech-language models (SLMs) offer a promising path toward unifying speech and text understanding and generation. However, challenges remain in achieving effective cross-modal alignment and high-quality speech generation. In this work, we systematically investigate the impact of key components (i.e., speech tokenizers, speech heads, and speaker modeling) on the performance of LLM-centric SLMs. We compare coupled, semi-decoupled, and fully decoupled speech tokenizers under a fair SLM framework and find that decoupled tokenization significantly improves alignment and synthesis quality. To address the information density mismatch between speech and text, we introduce multi-token prediction (MTP) into SLMs, enabling each hidden state to decode multiple speech tokens. This leads to up to 12 faster decoding and a substantial drop in word error rate (from 6.07 to 3.01). Furthermore, we propose a speaker-aware generation paradigm and introduce RoleTriviaQA, a large-scale role-playing knowledge QA benchmark with diverse speaker identities. Experiments demonstrate that our methods enhance both knowledge understanding and speaker consistency. 111Our demo, code, models, and the RoleTriviaQA dataset will be available at https://github.com/cnxupupup/SLM-Decoupled-MTP.
Speech-Language Models with Decoupled Tokenizers and
Multi-Token Prediction
Xiaoran Fan1††thanks: Equal contribution., Zhichao Sun111footnotemark: 1, Yangfan Gao111footnotemark: 1, Jingfei Xiong111footnotemark: 1, Hang Yan211footnotemark: 1, Yifei Cao1, Jiajun Sun1, Shuo Li1, Zhihao Zhang1, Zhiheng Xi1, Yuhao Zhou1, Senjie Jin1, Changhao Jiang1, Junjie Ye1, Ming Zhang1, Rui Zheng1, Zhenhua Han, Yunke Zhang3, Demei Yan3, Shaokang Dong3, Tao Ji1††thanks: Corresponding author., Tao Gui122footnotemark: 2, Qi Zhang122footnotemark: 2, Xuanjing Huang122footnotemark: 2 1Fudan University 2The Chinese University of Hong Kong 3Honor Device Co., Ltd {xrfan23, taoji, tgui, qz, xjhuang}@fudan.edu.cn
1 Introduction
In recent years, large language models (LLMs) have achieved remarkable progress in text understanding and generation Grattafiori et al. (2024); Qwen et al. (2025), fueling rapid advancements in multimodal models such as vision-language models Liu et al. (2023), vision-language-action models Kim et al. (2024), and speech-language models (SLMs) Chu et al. (2023); Li et al. (2025). Among them, SLMs are particularly valuable for applications like natural and fluent human-computer dialogue Défossez et al. (2024); Xu et al. (2025) and personalized speech generation Du et al. (2024a, c). The core challenge lies in generalizing the language understanding and generation capabilities of LLMs to incorporate and integrate speech understanding and generation.
Current approaches for connecting text and speech in SLMs can be broadly categorized into two perspectives. Speech-centric models are built upon mature speech generation architectures, with LLMs introduced as conditional context. However, while this approach achieves high-quality speech synthesis, it struggles to fully leverage LLMs’ rich world knowledge and powerful capabilities. LLM-centric models take LLMs as the backbone and extend them with speech interaction. In theory, it fully utilizes LLMs’ strengths (e.g., intent understanding, instruction following, planning, and decision-making). However, the quality of speech generation often suffers due to challenges in cross-modal alignment. This work focuses on the LLM-centric SLMs, aiming to explore how core components affect cross-modal alignment and the quality of generated speech.
The speech tokenizer and speech head determine the atomic inputs and latent representations of speech signals, which are critical to cross-modal alignment. Existing speech tokenizers can be categorized into three types: coupled tokenizers (e.g., WavTokenizer (2025)) that jointly encode semantic and acoustic details; semi-decoupled tokenizers (e.g., SpeechTokenizer (2024)) that distill semantic content via HuBERT (2021) while retaining partial coupling; and decoupled tokenizers (e.g., FACodec (2024)) that separate speech into independent subspaces such as semantics, prosody, and timbre. Although all three achieve comparable performance in speech reconstruction tasks, their compatibility with SLMs remains underexplored. Under fair comparison within SLMs, we find that decoupled tokenizers are more favorable for cross-modal alignment and yield higher speech synthesis quality.
We observe a mismatch in information density between speech and text representations in the speech head: one second of speech corresponds to hundreds of tokens, far exceeding the amount of text information humans typically express per second. In LLMs, multi-token prediction (MTP) is often used to accelerate inference or enrich hidden representations by enabling a single hidden state to predict multiple future tokens. To balance the information density between speech and text, we introduce MTP into SLMs, allowing each speech hidden vector to decode multiple speech tokens (for both coupled and decoupled tokenizers). We experimented with various compression ratios, ranging from one token per hidden state to twelve tokens per state. The MTP not only improved speech decoding speed by up to 12 but also significantly enhanced speech synthesis quality (the Word Error Rate decreased from 6.07 to 3.01).
In addition, we propose a speaker-aware speech generation paradigm, introducing speaker identity features into the speech-language context to guide and control the speaker timbre in synthesized speech. To comprehensively evaluate SLMs’ understanding and generalization of the general knowledge captured by LLMs, as well as their speaker consistency, we construct a role-playing knowledge QA task. Specifically, we use CosyVoice2 (2024b) to convert answers from a text-based knowledge QA dataset—TriviaQA (2017) into speech, assigning each answer to one of 15 Genshin characters. The resulting RoleTriviaQA dataset contains 138K/500/2,426 samples for training/validation/test, respectively. We reserve five characters and the WebQuestions (2025) test set for out-of-domain evaluation of both speaker consistency and knowledge QA performance (i.e., evaluating generalization). Results show that the decoupled architecture enables cooperative improvement in knowledge QA tasks and speaker timbre, and significantly outperforms other baselines.
2 Speech-Language Model
A typical speech-language model (SLM) is usually an extension of a successful large language model (LLM) capable of understanding and generating both text and speech signals. It relies on a speech-language tokenizer to convert raw text or audio into discrete tokens, a decoder-only Transformer to learn cross-modal context and alignment, a language head to predict text, and a speech head to predict speech tokens then synthesize continuous audio signals.
In the following, we introduce these core components and discuss how different design choices affect the SLM’s capabilities, particularly in the speech-language tokenizer and speech head.
2.1 Core Components
Speech-Language Tokenizer
is a key component of SLMs, as it defines the atomic representations of text and speech. High-quality representations are crucial for effective cross-modal alignment, efficient generation, and preservation of acoustic details, making this component the focus of extensive research Zhan et al. (2024); Peng et al. (2024). It converts text or speech into discrete token sequences, such as for text and for speech. These sequences are ultimately combined into interleaved text-speech sequences to handle a variety of speech-language tasks, such as ASR Fathullah et al. (2024), TTS Du et al. (2024a, b), speech translation Chen et al. (2024b), and spoken question answering Zhao et al. (2024).
Speech tokenizers can be categorized into coupled and decoupled 222 For simplicity, we assume a two-way decomposition, though in practice it can be readily extended to more components, e.g., . , 333 Decoupled sequences may also be arranged in a type-wise contiguous manner within a chunk (e.g., ) rather than in an interleaved fashion. However, experiments in Appendix 8 show that interleaved ordering yields significantly better performance. designs for semantic and acoustic details . Intuitively, as illustrated in Figure 1, decoupled semantic tokens are easier to align with text tokens, while the SLM can leverage its overparameterization to handle the acoustic detail tokens specifically. In contrast, coupled tokenizers may compromise cross-modal semantic alignment.
However, current comparisons between speech tokenizers are primarily conducted on speech reconstruction tasks, which are independent of SLMs. As a result, it is difficult to accurately assess their suitability for SLMs. For example, under fair conditions, the coupled BigCodec (2024) achieves better reconstruction quality, yet when integrated into an SLM, the decoupled FACodec (2024) yields significantly better performance. This paper provides a detailed analysis of mainstream speech tokenizers within the context of SLM evaluation.
Decoder-only Transformer
has achieved great success as a foundation model for text understanding and generation. After unifying the tokenization of text and speech, it can further learn joint speech-text context and perform cross-modal alignment. Its parameters are typically initialized from a well-pretrained LLM and then adapted to multimodal contexts. Given a sequence of input tokens , it produces a hidden representation vector for each token,
Prediction Heads
include a language head and a speech head. The language head is consistent with that of an LLM and can be initialized from pretrained parameters. In contrast, the speech head has a different “vocabulary size” and is randomly initialized. Language head consists of a linear transformation followed by a softmax function,
A parameter matrix maps the hidden vector to the vocabulary space of size . During training, the model is optimized using the cross-entropy loss: , and during inference, the next token is predicted by either taking the or sampling.
Speech head follows a similar formal definition, training objective, and inference process as the language head,
It uses a parameter matrix to map the hidden vector to the speech “vocabulary” of size . When semantic and acoustic components are decoupled (i.e., is split into and ), the vocabulary and the linear projection are also split accordingly into independent parts:
where and .
Finally, the speech is reconstructed through the audio decoder from a pretrained codec:
2.2 Multi Token Prediction
We observe that text representations are significantly more information-dense than speech. Typically, one second of speech is decoded into several hundred tokens, while it rarely conveys more than ~20 text tokens’ worth of information. This asymmetry in information density can make cross-modal alignment difficult to learn. On the other hand, speech tokens with excessively low information density can also hinder SLM inference efficiency, as generating one second of speech may take significantly longer than one second. Therefore, balancing the information density between speech and text representations becomes a critical challenge.
Multi-token prediction (MTP) Gloeckle et al. (2024); Cai et al. (2024); Liu et al. (2024) is commonly used in LLMs to accelerate inference or enrich hidden representations (e.g., by requiring to decode not only but also future tokens like ). To compress the information density of speech representations, we introduce MTP into the SLM framework. As depicted in Figure 2, compared with Next Token Prediction (NTP), MTP enables each hidden vector to decode multiple speech tokens, whether coupled or decoupled, thus alleviates the imbalance in information density between speech and text.
We group every adjacent speech tokens into a multi-token group , which is decoded simultaneously at step , meaning all tokens in are predicted from the same hidden vector . This significantly increases the information density encoded in , making it more compatible with text tokens. The definitions of the multi-token groups / 444 Where the group consists of half semantic tokens and half acoustic detail tokens . for coupled/decoupled tokenizers subject to are as follows:
The parameter matrix of the speech head is extended from a 2D tensor to a 3D tensor to enable simultaneous prediction of multiple tokens. The extended dimension has size , where each slice corresponds to a linear projection matrix (as in Equation 2 for coupled or Equation 3 for decoupled tokens) for one token in the group ,
Similarly, the softmax normalization is applied after the linear transformations to yield the probability distributions for all tokens in the group,
The loss / is also extended to the average over all tokens within the group,
Finally, the Transformer input must be adapted for MTP. For each group, we first obtain the embeddings of all tokens (i.e., in NTP), concatenate them, and feed the result into a fusion network like a linear or MLP to produce a downsampled (by a factor of ) input vector:
where the is the concatenate function.
2.3 Speaker-Aware Speech Generation
In speech generation (especially within spoken dialogue systems), explicitly controlling paralinguistic features such as speaker timbre, intonation, and emotional tone enables the generation of consistent and controllable speech responses. Therefore, we introduce speaker-specific representations into the SLM context and evaluate both speech quality and character consistency in a role-playing speech QA task.
Inspired by Du et al. (2024a) and Li et al. (2024), we employ a pretrained timbre extractor 555 We adopt the timbre extractor from the decoupled tokenizer FACodec Ju et al. (2024) in this work. Implementation details are provided in Appendix B. It can be smoothly replaced with other extractors if needed. to obtain the timbre representation of the speaker or role. This representation can either be a discretized token , or a continuous vector sequence , where is the sequence length. The speech-language context of the SLM is then extended as follows:
We expect the generated speech to be guided and controlled by the speaker-specific features.
3 Role-Playing Knowledge QA Task
To evaluate which SLM architecture better aligns and generalizes the inherent world knowledge embedded in LLMs, and which design better preserves speaker or character consistency, we construct a role-playing knowledge QA task. On the data side, we introduce RoleTriviaQA, a role-playing speech knowledge QA dataset. On the model training side, we propose a two-stage training framework. An overview of our proposed model is illustrated in Figure 3.
3.1 RoleTriviaQA
Inspired by the text knowledge QA— TriviaQA (2017), we introduce a role-playing speech QA extension. It requires the SLM to retain the LLM’s knowledge and then generate accurate speech consistent with a specific role. To build a high-quality role-playing knowledge QA dataset, we design a two-stage data construction pipeline: (1) Speaker selection from Genshin Impact: 10 seen speakers for both training and testing, and 5 unseen only for testing; (2) Speech synthesis via CosyVoice 2 (2024b), leveraging filtered high-quality reference speeches to synthesize the role-specific speech of the text answers.
This pipeline yields 138,384 training samples, with 1,500 in-domain and 926 out-of-domain evaluation samples. More details of RoleTriviaQA are provided in Appendix G.
Model | Codebook Size | Nq | Token Rate (TPS) | Success Rate (SR) | UTMOS | WER | SIM | |
Seen | Unseen | |||||||
Coupled Tokenizers trained with SLM | ||||||||
Encodec (2023) | 1024 | 4 | 300 | 0.38 | 2.86 | 17.00 | 0.18 | 0.10 |
Encodec (2023) | 1024 | 8 | 600 | 0.24 | 3.50 | 8.14 | 0.18 | 0.10 |
WavTokenizer (2025) | 4096 | 1 | 75 | 0.38 | 2.80 | 89.60 | 0.06 | 0.04 |
WavTokenizer-v2 (2025) | 4096 | 1 | 75 | 0.65 | 3.16 | 49.24 | 0.11 | 0.08 |
StableCodec (2024) | 15625 | 2 | 50 | 0.98 | 3.90 | 13.04 | 0.17 | 0.14 |
BigCodec (2024) | 8192 | 1 | 80 | 0.70 | 3.93 | 13.63 | 0.18 | 0.10 |
Semi-Decoupled Tokenizers trained with SLM | ||||||||
SpeechTokenizer (2024) | 1024 | 5 | 250 | 0.99 | 3.86 | 13.13 | 0.16 | 0.13 |
SpeechTokenizer (2024) | 1024 | 8 | 400 | 0.90 | 4.05 | 9.95 | 0.14 | 0.14 |
Decoupled Tokenizers trained with SLM | ||||||||
FACodec-NTP-1H | 1024 | 3 | 240 | 0.87 | 3.93 | 6.07 | 0.50 | 0.49 |
FACodec-MTP-3H | 1024 | 3 | 80 | 0.82 | 3.83 | 5.84 | 0.47 | 0.47 |
FACodec-MTP-6H | 1024 | 3 | 40 | 0.92 | 3.75 | 4.36 | 0.47 | 0.47 |
FACodec-MTP-12H | 1024 | 3 | 20 | 1.00 | 3.67 | 3.01 | 0.47 | 0.47 |
3.2 Training
To build an SLM support role-playing knowledge QA task, we adopt a two-stage training method:
Stage 1: Speech-Language Pretraining
To improve cross-modal alignment, we conduct a large-scale dataset Emilia-3.5. First, we filter Emilia (2024) (a large-scale, real-world speech collection with diverse emotions and styles) for samples scoring at least 3.5 on DNSMOS (2022). Next, we merge the standard LibriTTS (2019) training set. The resulting corpus comprises 2,495 hours of high-quality speech.
Moreover, to enable bidirectional alignment, we conduct both ASR and TTS training. Specifically, we adopt as the TTS data format and for ASR. The pretaining loss is formulated as:
where the predicted token could be a single token or a multi-token group (i.e., ).
Stage 2: Instruction Fine-Tuning
After the pretraining stage, we perform supervised fine-tuning (SFT) with RoleTriviaQA. The data format is , and the SFT loss is defined as follows:
where the loss is calculated across both the text answer and the speech answer.
4 Experiment
Our experiments address two critical questions:
- 1.
2.
How can SLMs effectively leverage their internal knowledge while maintaining high-quality speech synthesis? (Section 4.4)
Setups
Details of the datasets, baselines, evaluation, and hyperparameter setups are placed in Appendix C.
4.1 Effect of Speech Tokenizers
As shown in Table 1, the degree of speech representation decoupling shows a significant correlation with the joint SLM training effect, where the semi-decoupled and fully decoupled architectures exhibit better overall performance.
First, coupled architectures exhibit severe convergence challenges during SLM training (shown in Fig. 4), resulting in generally low success rates of speech synthesis (70%). The notable exception is StableCodec, whose large codebook (15,625 entries) and low token rate (50 TPS) establish dual compensation mechanisms that mitigate inherent limitations of coupled architectures. Second, the semi-decoupled and decoupled architectures show clear advantages. The semi-decoupled architecture (SpeechTokenizer) achieves the best speech quality (UTMOS), while the decoupled architecture works optimally in WER and SIM. Third, the decoupled architectures’ SIM between the training and test sets is consistently larger and significantly better than the other baselines, suggesting superior generalization capabilities, especially for speaker timbre.
Overall, decoupled tokenizer and SLM joint modeling is better, and has better controllability and scalability.
4.2 Next Token vs. Multi-Token Prediction
As illustrated in Table 1, to investigate the effectiveness of MTP under decoupled architecture, we also conduct experiments comparing models trained with different compression ratios.
First, the MTP-3H model is competitive with the NTP-1H, even with compression. Moreover, it has also surpassed all of the semi-decoupled and coupled baselines in WER and SIM. Second, when increasing compression from (MTP-3H) to (MTP-12H), WER improves by 48% (5.84 to 3.01), with stable SIM and UTMOS experiencing a slight decline. Notably, under 12 compression, our model achieves speech synthesis with a 100% success rate, and the WER (3.01) is comparable to the ground truth. These results reveal that speech length compression based on the decoupled tokenizer effectively mitigates the frequency mismatch between text (~20Hz) and speech (240Hz) modalities, enhancing cross-modal alignment performance.
4.3 Speaker-Aware Speech Generation
Model | H | TPS | Spk-Aware | UTMOS | WER | SIM |
NTP | 1 | 240 | ✗ | 3.93 | 6.07 | 0.49 |
NTP | 1 | 240 | ✓ | 4.05 | 5.51 | 0.57 |
MTP | 3 | 80 | ✗ | 3.83 | 5.84 | 0.47 |
MTP | 3 | 80 | ✓ | 3.90 | 7.04 | 0.56 |
MTP | 6 | 40 | ✗ | 3.75 | 4.36 | 0.47 |
MTP | 6 | 40 | ✓ | 3.94 | 4.07 | 0.59 |
MTP | 12 | 20 | ✗ | 3.67 | 3.01 | 0.47 |
MTP | 12 | 20 | ✓ | 3.93 | 2.94 | 0.60 |
Models | Model Type | Tokenizer Type | In-Domain (RoleTriviaQA) | Out-of-Domain (Web Questions) | ||||||||
EM(%) | F1(%) | SIM | EM | F1 | SIM | |||||||
Seen | Unseen | All | Seen | Unseen | All | |||||||
Qwen2.5-0.5B-Instruct | LLM | - | 7.1 | 17.3 | - | - | - | 1.1 | 14.0 | - | - | - |
WavTokenizer-v2 | SLM | Coupled | 5.1 | 12.5 | 0.22 | 0.18 | 0.20 | 1.1 | 3.9 | 0.21 | 0.18 | 0.20 |
BigCodec | SLM | Coupled | 2.9 | 9.4 | 0.23 | 0.17 | 0.21 | 1.0 | 4.2 | 0.23 | 0.17 | 0.21 |
StableCodec | SLM | Coupled | 7.5 | 13.8 | 0.18 | 0.14 | 0.17 | 0.6 | 3.6 | 0.17 | 0.14 | 0.16 |
SpeechTokenizer | SLM | Semi-Decoupled | 3.4 | 9.1 | 0.19 | 0.14 | 0.17 | 0.3 | 2.6 | 0.20 | 0.14 | 0.18 |
FACodec | SLM | Decoupled | 12.0 | 23.8 | 0.63 | 0.58 | 0.61 | 9.0 | 17.9 | 0.63 | 0.58 | 0.61 |
Table 2 presents that Speaker-Aware training exhibits consistent enhancement effects across NTP and MTP architectures with decoupled tokenizer.
First, for NTP, the incorporation of speaker embedding leads to consistent improvements across all metrics. Second, for MTP, speaker-embedding provides a stable gain in most cases. In addition, as compression increases from 3H to 12H, Speaker-Aware models also show continuous WER reduction (7.04 to 2.94), In particular, the advantages of Speaker-Aware training become more evident at higher compression. For instance, the MTP-12H trained with speaker embedding significantly outperforms its non-aware model.
This indicates that incorporating speaker embeddings enables the model to better capture speaker’s paralinguistic features and linguistic accuracy.
4.4 Role-Playing Knowledge QA
Settings
To evaluate the capability of our decoupled SLM in preserving both knowledge and acoustic features, we compare it against the top-3 coupled and semi-decoupled SLM baselines (based on Table 1) and the LLM backbone Qwen2.5-0.5B-Instruct.
Results
Table 3 shows that the decoupled model establishes state-of-the-art performance across all metrics. First, for in-domain evaluation, our decoupled model achieves 12.0 EM and 23.8 F1, surpassing all baselines, while its 0.61 SIM score is 2.9 higher than the best coupled baseline (BigCodec). For out-of-domain evaluation, while all models degrade, ours exhibits superior robustness. This highlights that decoupling enables joint optimization of semantic and acoustic objectives, whereas semi-decoupled/coupled systems suffer from target conflicts. Second, compared with the LLM baseline, our decoupled SLM significantly outperforms on both ID and OOD evaluations. Notably, for OOD, the EM (9.0 vs 1.1) and the F1 (17.9 vs 14.0), demonstrating its effectiveness in preserving LLM knowledge while adapting to unseen queries. Particularly, for SIM generalization, our model achieves high SIM scores on both seen/unseen speakers (0.63 vs 0.58), with minimal performance gaps. This demonstrates its strong generalization in preserving speaker timbre.
Overall, our results confirm that the decoupled architecture enables cooperative improvement in knowledge QA tasks and speaker timbre, and outperforms other baselines.
Models | Heads | TPS | Cosine Similarity | Euclidean Distance | Riemannian Distance | ||||
TTSim | SSSim | STSim | TTDist | SSDist | STDist | ||||
NTP | 1 | 240 | 0.896 | 0.999 | 0.480 | 0.438 | 0.030 | 1.014 | 365.26 |
MTP | 3 | 80 | 0.986 | 0.999 | 0.967 | 0.147 | 0.052 | 0.244 | 288.39 |
MTP | 6 | 40 | 0.991 | 0.998 | 0.981 | 0.116 | 0.055 | 0.186 | 268.19 |
MTP | 12 | 20 | 0.996 | 0.998 | 0.986 | 0.078 | 0.063 | 0.158 | 237.55 |
4.5 Ablation Study
MTP’s Architecture Selection We analyze fusion strategies and speech head architectures to select the optimal MTP configuration, as summarized in Table 6. NTP’s Layer and Token Organization Selection We explore the best number of decoupled tokenizer codebook layers and speech token organization patterns for SLM training. Details can be found in Section E.2 and Table 8.
4.6 Effect of MTP on Cross-modal Alignment
Table 1 and Table 2 shows that higher speech token compression in the MTP architecture consistently improves performance, especially in WER. To investigate the hypothesis that higher compression improves modal alignment, we conduct detailed quantitative and qualitative analysis.
Quantitative Analysis
As shown in Table 4, token compression improves cross-modal alignment and maintains intra-modal integrity, with MTP models outperforming NTP baselines. First, when compression increases from 3× to 12×, intra-modal similarities (TTSim, SSSim) remain high in all settings, while cross-modal similarity (STSim) increases steadily from 0.967 to 0.986. Second, the cross-model distance (STDist) decreases significantly from 0.244 to 0.158 , while intra-modal distances (TTDist, SSDist) stay stable. These trends align with WER improvements, suggesting that stronger compression encourages more isomorphic and effective speech-text representations. In addition, compared to NTP, all MTP settings show better cross-modal alignment. This demonstrates that MTP not only maintains intra-modal coherence but also enhances cross-modal correspondence under token compression.
Moreover, we use Riemannian distance (2009) to quantify speech-text alignment, capturing the intrinsic geometry of the embedding space by modeling it as a curved manifold. The distance between the dominant speech hidden states and text hidden states is computed as:
where is the -th positive real eigenvalue of , and are the centroids of speech and text embeddings respectively. Here can be calculated from the SVD of the right singular matrices : With higher compression, the Riemannian distance decreases (288.39 to 237.55), indicating that aggressive compression not only reduces embedding distance but also better preserves their intrinsic geometric structure.
Qualitative Analysis
We include complementary qualitative results in Section F.1.
4.7 Effect of Tokenizer-MTP Combination
To examine the effectiveness of different tokenizer coupling strategies in MTP SLMs, we perform comprehensive experiments in Section F.2.
5 Related Work
Speech Tokenization
have advanced through RVQ-based architectures and quantization innovations: SoundStream (2022) introduced the RVQ-VAE framework, enhanced by Encodec (2023, LSTM temporal modeling) and SpeechTokenizer (2024, HuBERT semantic distillation). Recent ultra-low-bitrate methods include Improved RVQGAN (DAC, 2023) with quantizer dropout, WavTokenizer (2025, 0.9kbps via single-layer RVQ), BigCodec (2024, 1.04kbps via Convolution-LSTM projection), and StableCodec (2024, 0.7kbps using Transformer with FSQ (2024)).
Speech Language Models
LLMs are increasingly adopted for unified speech processing through two primary paradigms. The first integrates speech encoders with LLMs for comprehension tasks like ASR Ma et al. (2024) and spoken language understanding Chu et al. (2023), achieving strong analytical capabilities but lacking generative functionality. The second targets speech generation via dual strategies: codec-based models (e.g., SpeechGPT, 2023) discretize speech into compressed tokens, while diffusion-enhanced approaches (e.g., CosyVoice, 2024a; 2024c) synthesize waveforms through LLM-diffusion.
Multi Token Prediction
has emerged as a critical technique in LLMs to accelerate inference. Detailed related works are placed on Appendix H.
6 Conclusion
This work investigates how LLM-centric SLMs can better align and generalize speech-text understanding and generation by revisiting core design choices. We find that fully decoupled tokenizers and multi-token prediction significantly improve alignment efficiency and speech quality. Additionally, incorporating speaker-aware modeling and role-based QA evaluation reveals improved generalization and character consistency. Our findings provide practical insights for building more capable and scalable SLMs.
Limitations
Verification on Other Decoupled Speech Tokenizers
While our decoupled SLM framework shows promising results, broader validation across comparable systems remains advisable. The current evaluation is limited to FACodec as no more open-source public implementations can be found in this emerging field to the best of our knowledge. Future work can extend validation to other decoupled speech tokenizers to better assess generalizability of our method.
Challenges under High Compression
Although MTP consistently improves cross-modal alignment and reduces the Word Error Rate (WER), achieving near ground truth, even at 12× compression. However, this aggressive compression results in a slight decline in UTMOS. This quality degradation probably stems from the loss of fine-grained acoustic details at a high compression ratio. Our future work will investigate adaptive compression strategies or richer intermediate representations to better enhance the synthesis quality.
Appendix A Details of Baselines
A.1 Compared Baselines
Encodec Défossez et al. (2023): An RVQ-based codec designed for universal audio compression.666https://huggingface.co/facebook/encodec_24khz
WavTokenizer Ji et al. (2025): A single VQ codebook-based tokenizer for universal audio.777https://huggingface.co/novateur/WavTokenizer-large-speech-75token
StableCodec Parker et al. (2024): A residual FSQ-based tokenizer for speech.888https://huggingface.co/stabilityai/stable-codec-speech-16k
BigCodec Xin et al. (2024): A VQ-based single-stream codec for speech.999https://huggingface.co/Alethia/BigCodec/resolve/main/bigcodec.pt
SpeechTokenizer Zhang et al. (2024): An RVQ-based codec with semantic distillation for speech.101010https://huggingface.co/fnlp/AnyGPT-speech-modules/tree/main/speechtokenizer
A.2 Results of Codec Reconstruction
Model |
Nq |
STOI |
UTMOS | SIM | |||||||
Ground Truth | - | - | - | - | 1.00 | 4.55 | 4.09 | 1.00 | |||||||
Encodec | 1024 | 8 | 600 | 6000 | 0.94 | 3.18 | 3.09 | 0.88 | |||||||
WavTokenizer | 4096 | 1 | 75 | 900 | 0.90 | 2.63 | 3.79 | 0.65 | |||||||
StableCodec | 15625 | 2 | 50 | 697 | 0.91 | 2.91 | 4.23 | 0.62 | |||||||
BigCodec | 8192 | 1 | 80 | 1040 | 0.93 | 3.27 | 4.11 | 0.84 | |||||||
SpeechTokenizer | 1024 | 8 | 100 | 1000 | 0.92 | 2.61 | 3.87 | 0.83 | |||||||
FACodec | 1024 | 6 | 480 | 4800 | 0.95 | 3.47 | 3.90 | 0.87 |
Table 5 shows the results of VQ reconstruction evaluation on the LibriTTS test-clean dataset. The results reveal that for acoustic quality (STOI Taal et al. (2010), PESQ Rix et al. (2001) and UTMOS), all tokenizers perform similar with the ground truth, sharing close metric values. But for SIM, the trend differs: the VQ performs better on SIM as its token rate increases, and the decoupled model (FACodec) surpasses most of the coupled and semi-decoupled models. This is probably because reduced token rates compromise timbre fidelity, combined with coupled modeling diminishing acoustic prominence relative to semantic components.
Appendix B Speaker Embedding Extraction
In FACodec, speaker information is extracted by a timbre extractor which is composed of 4 transformer encoder blocks. The module operates on the output of the speech encoder and transforms it into a global vector , which captures the characteristics of speakers. To generate audios, we sum the representations of prosody, content, and then integrate the speaker embedding via conditional layer normalization, enabling effective conditioning on speaker timbre during generation.
Appendix C Experimental Setups
Datasets
For the default TTS training, we utilized the LibriTTS dataset, which contains 585 hours of English speech data. All audio samples are resampled to 16 kHz.
For Role-Playing Knowledge QA, we leveraged the proposed Emilia-3.5 (2,495 hours) and RoleTriviaQA (138k) datasets during stage 1 and stage 2. Our proposed data details can be found in Section 3.
Baselines
We selected the current state-of-the-art speech tokenizer as the baseline for our model. To ensure a fair comparison, we employed the official weight files provided by the Encodec Défossez et al. (2023), SpeechTokenizer Zhang et al. (2024), WavTokenizer Ji et al. (2025), StableCodec Parker et al. (2024), BigCodec Xin et al. (2024). Details can be found in Section A.1.
Training and Inference Hyperparameters
Our baseline training configuration employs the AdamW optimizer ( and ) with a 5e-4 initial learning rate following a cosine decay schedule. All models undergo 90k training steps on 8 A800 GPUs using consistent batch sizes of 16 for fair comparison. For specialized tasks like Role-Playing Knowledge QA, we implement a two-phase adaptation: 40k steps for cross-modal alignment pretraining followed by 12k steps of knowledge-role joint fine-tuning. During inference, we apply greedy decoding with a 1.2 repetition penalty to balance output quality and diversity.
Training Loss Comparison
Figure 4 depicts the training loss of NTP TTS training with LibriTTS and 2-stage training of Role-Playing Knowledge QA task of all coupled, semi-decoupled baselines and FACodec, on which tokenizer types are annotated. It can be clearly seen that decoupled and semi-decoupled SLMs perform similarly and significantly better than coupled baselines across all tasks and stages.
Evaluation Setups
For the default TTS evaluation, following Zhang et al. (2024), we randomly sample 300 samples from the LibriTTS test set for speech synthesis evaluation. For Role-Playing Knowledge QA, the evaluation set of RoleTriviaQA contains 1,500 in-domain and 926 out-of-domain samples.
Metrics
For the default TTS evaluation, we evaluate Success Rate (SR), speech quality (UTMOS Saeki et al. (2022)), Word Error Rate (WER), and speaker similarity (SIM).
For Role-Playing Knowledge QA task, we report three metrics: Exact Match (EM) and F1 score for text responses, while SIM for speech responses. Details of these metrics can be found in Appendix D.
Appendix D Details of Evaluation Metrics
For SLM performance evaluation, we adopted:
Success Rate (SR) measures the model’s ability to learn token sequence patterns. It is defined as the ratio of successfully synthesized speech samples to the total number of test samples:
Word Error Rate (WER) measures the accuracy of the semantic information of the generated speech. We employed the ASR model Whisper-large-v3111111https://huggingface.co/openai/whisper-large-v3 to transcribe generated speech into text and then calculate the WER between the transcribed text and the text response.
UTMOS assesses the acoustic quality of the generated speech. We utilized the model provided by SaruLab121212https://huggingface.co/spaces/sarulab-speech/UTMOS-demo/resolve/main/epoch%3D3-step%3D7459.ckpt to evaluate it.
Speaker Similarity (SIM) reflects how similar it is between the generated speech and the reference speech. We considered the WavLM-TDCNN Chen et al. (2022) speaker embedding model to assess speaker similarity.
2 more metrics are considered for the evaluation of the VQ reconstruction (See Appendix A.2):
PESQ quantifies the perceptual quality of speech signals by simulating human auditory perception under various distortions and degradations.
STOI predicts speech intelligibility by analyzing time-frequency correlations between clean and processed speech signals.
For text quality evaluation, we utilized:
Exact Match (EM) is the percentage of generated answers that exactly match the ground truth, which is given by:
where is the number of samples, is the generated answer, and is the ground truth, with representing the indicator function.
F1 score represents the harmonic mean of precision and recall, formulated as:
Precision Recall F1 where TP denotes words appearing in both ground truth and predicted answers, FP those exclusively in predictions, and FN those solely in ground truth.
Appendix E Ablation Study
E.1 Effect of MTP Architecture
Fusion | Speech Head | UTMOS | WER | SIM |
Linear | Linear | 3.75 | 9.42 | 0.45 |
Linear | MLP | 3.78 | 7.28 | 0.46 |
Linear | Transformers | 3.31 | 22.19 | 0.41 |
MLP | MLP | 3.75 | 11.8 | 0.45 |
MLP | Transformers | 3.34 | 16.84 | 0.41 |
MLP | Linear | 3.90 | 6.35 | 0.47 |
Table 6 compares different MTP architectures (Fusion and Speech Head). First, controlling the speech head, compared to Linear, the MLP fusion improves performance in most cases (2/3). The results demonstrate the importance of non-linear fusion to decouple speech features. Transformer-based heads result in substantial degradation in the WER, regardless of the fusion method, likely due to overfitting or incompatibility with the fused representations. The MLP–Linear configuration achieves the strongest overall performance in all metrics. Therefore, we select it as the default experimental setting for MTP.
E.2 Effect of Number of Codebook Layers in SLMs
Nq |
UTMOS | WER | SIM | ||||
3 | FWI | 0.87 | 3.93 | 6.07 | 0.49 | ||||
6 | FWI | 0.82 | 4.00 | 8.23 | 0.34 |
As indicated in Table 7, the 3-layer setting (1 prosody layer and 2 content layers) in FACodec-NTP training exhibits superior overall performance and higher success rate compared to the full 6-layer (adding 3 acoustic layers based on the previous setting) configuration. Therefore, we chose the 3-layer setting for main NTP and MTP experiments.
E.3 Effect of Speech Token Organization Pattern
Nq |
UTMOS | WER | SIM | ||||
3 | CWI | 0.80 | 3.30 | 22.21 | 0.43 | ||||
3 | FWI | 0.87 | 3.93 | 6.07 | 0.49 |
We compared two speech token organization patterns: Frame-Wise Interleaving (FWI) and Chunk-Wise Interleaving (CWI).
FWI
organizes tokens as per frame, where denotes prosody tokens and represents content tokens for the -th frame.
CWI
groups tokens into chunks with layer length (typically 80), where the final chunk may have 80 tokens.
Table 8 shows that FWI outperforms CWI on all metrics especially on WER (6.07 vs. 22.21), which means the CWI-trained model aligns text and speech modalities less effectively than its FWI-trained counterpart. Therefore, we opt FWI as the default setting for our main experiments.
Appendix F Analysis
F.1 Quantitative Analysis of MTP’s Impact on Cross-Modal Alignment
Figure 5 displays UMAP projections of speech and text representations from different layers in SLM. These projections illustrate the impact of different compression rates on the SLM’s training, revealing three key observations.
First, speech and text representations consistently form distinct modality-specific clusters across all layers and compression rates. Second, all models show a layer-wise refinement. Although cross-modal distances increase in intermediate layers, deeper layers progressively reduce the gap between modalities, where cross-modal distances are lower in the final layer compared to the word embedding layer. Third, higher compression consistently reduces cross-modal distance. Specifically, 12H shows the closest alignment in both word embedding (), mid layers (), and last layers (), followed by 6H and 3H, indicating that compression facilitates tighter alignment between modalities.
F.2 Effect of Tokenizer-MTP Combination
Models | CR | UTMOS | WER | SIM |
Coupled Tokenizers trained with SLM | ||||
BigCodec | 1 | 3.93 | 13.63 | - |
BigCodec | 2 | 3.97 | 15.48 | - |
BigCodec | 4 | 3.29 | 13.48 | - |
BigCodec | 8 | 1.94 | 46.82 | - |
Semi-Decoupled Tokenizers trained with SLM | ||||
SpeechTokenizer | 1 | 3.86 | 13.13 | - |
SpeechTokenizer | 5 | 3.57 | 12.9 | - |
SpeechTokenizer | 10 | 3.23 | 9.67 | - |
Decoupled Tokenizers trained with SLM | ||||
FACodec | 1 | 3.93 | 6.07 | 0.49 |
FACodec | 3 | 3.90 | 6.35 | 0.47 |
FACodec | 6 | 3.75 | 4.36 | 0.47 |
FACodec | 12 | 3.67 | 3.01 | 0.47 |
Table 9 investigates the impact of different tokenizers trained on MTP based SLMs. We select semi-decoupled (SpeechTokenizer), decoupled (FACodec) tokenizers for comparsion, and choose BigCodec as our coupled tokenizer baseline, due to its 80 TPS for fair analysis and strong performance in coupled tokenizers as shown in Table 1.
Primarily, coupled models experience severe degradation as compression increases: UTMOS drops sharply from 3.97 to 1.94, and WER rises to 46.82 at 8× compression. Additionally, semi-decoupled models show some WER improvement but with a slight decline in UTMOS, even with 10x compression rate. Finally, the decoupled architecture maintains speaker similarity, and consistently reduces the WER under higher compression, with only a minor reduction in audio quality. Remarkably, even at a 12 compression, the WER is close to ground truth. Our above results indicate that: MTP architectures achieve the best performance with decoupled tokenizers, followed by semi-decoupled and coupled tokenizers.
Appendix G Details of RoleTriviaQA
Speaker | Gender | Train Dataset Size | Evaluation Dataset Size | |
TriviaQA (ID) | Web Questions (OOD) | |||
Seen speakers | ||||
Madame Ping | Old Female | 14,244 | 100 | 65 |
Venti | Teenage Male | 13,664 | 100 | 47 |
Xiao | Male | 14,111 | 100 | 54 |
Kaedehara Kazuha | Male | 13,570 | 100 | 55 |
Arataki Itto | Male | 13,751 | 100 | 58 |
Kamizato Ayato | Male | 14,025 | 100 | 64 |
Furina | Female | 14,266 | 100 | 58 |
Nahida | Female | 13,357 | 100 | 69 |
Yae Miko | Female | 13,807 | 100 | 73 |
Keqing | Female | 13,589 | 100 | 80 |
Unseen speakers | ||||
Albedo | Male | - | 100 | 45 |
Wriothesley | Male | - | 100 | 70 |
Jean | Female | - | 100 | 64 |
Lisa | Female | - | 100 | 62 |
Eula | Female | - | 100 | 62 |
Total | 138,384 | 1,500 | 926 |
G.1 Data Construction
Data Composition
Table 10 demonstrates the detailed composition of RoleTriviaQA along with the selected speakers, with the ID evaluation samples are collected from TriviaQA test set and OOD samples are gathered from Web Questions of OpenAudioBench Li et al. (2025). Notably, we added about 26k text QA training samples filtered from LLaVA-Instruct-150k131313https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K during training to preserve the original textual ability of the SLM.
Role-Playing Speaker Selection
For seen speakers, we selected 10 representative speakers, including 5 males and 5 females, from 130 common speakers based on the character’s hotness ranking and the quality of synthesized speeches while maintaining age diversity. For unseen speakers, we selected 5 speakers, including 2 males and 3 females, from the remaining 120 speakers with the same criteria.
Speech Synthesis
For each character’s speech synthesis, we select high-quality reference speeches by combining speaker embedding similarity metrics with manual filtering as an comprehensive criteria. Then we leverage the CosyVoice 2 141414https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B system to generate role-playing speech from text answers. Finally, all speeches are resampled to 16 kHz to fit the tokenization requirement.
G.2 Samples of RoleTriviaQA
G.3 Prompt template used in Knowledge-Role Joint Fine-Tuning
Appendix H Multi Token Prediction
Multi Token Prediction (MTP) is leveraged in LLMs to accelerate inference through three principal strategies: 1) Multi-head architectures enhance generation quality via parallel output heads Gloeckle et al. (2024); 2) Tree-based systems like MEDUSA Cai et al. (2024) refine candidate selection through hierarchical validation; 3) Scalable frameworks such as DeepSeek-V3 DeepSeek-AI et al. (2024) enable efficient large-scale training. Recent adaptations extend MTP to speech processing: SLAM-Omni Chen et al. (2024a) employs audio token grouping for computational efficiency but neglects paralinguistic modeling through semantic-focused processing, while VocalNet Wang et al. (2025) achieves localized acceleration via DeepSeek-V3-style parallel prediction yet faces pipeline complexity from flow-matching requirements.